# Enhancing Supervised Composed Image Retrieval via Reasoning-Augmented Representation Engineering

### v1

@@ Introduction
In recent years, Composed Image Retrieval (CIR) (Vo et al. 2019a; Goenka et al. 2022; Wen et al. 2023; Bai et al. 2024) has gained increasing attention. Compared to traditional image retrieval tasks (Gordo et al. 2016; Liu et al. 2016; Chen et al. 2022), CIR requires not only a reference image as input but also a textual description of the expected modifications to the reference image. The complexity of understanding multi-modal data and modeling cross-modal relationships presents technical challenges, but also enables novel applications of CIR methods in fields such as e-commerce (Shankar et al. 2017) and Internet search (Deldjoo et al. 2023).
@@
Large Vision-Language Models (LVLMs) have demonstrated promising multimodal understanding capabilities, and have been adopted in various multimodal tasks, including visual grounding (Deng et al. 2021), visual question answering (Antol et al. 2015), and CIR tasks (Bai et al. 2024; Sun et al. 2025; Karthik et al. 2024; Sun, Ye, and Gong 2023). However, some of these works (Karthik et al. 2024; Tang et al. 2025b) use LVLMs to extract better textual descriptions, which are then used to retrieve images, while others (Sun et al. 2025; Bai et al. 2024) rely on complex input prompt designs. These cross-modal generation—such as image captioning—inevitably leads to information loss. For that sake, these methods may fall short in CIR tasks which require both multi-granular visual information and cross-modal correlation understanding. For example, given the modified instruction ‘dog sticks out its tongue on the grass’ (in Fig. 1), the CIR models need to consider both macro-level background information ‘on the grass’ and micro-level details of the dog’s tongue.
@@
Based on these considerations, we propose training a Pyramid Matching Model by utilizing a pre-trained LVLM. By removing the large language model (LLM) head and taking the last token as representation, we simply input composed queries and target images, and use InfoNCE (He et al. 2020) loss to minimize the distance between positive pairs while maximizing the distance between negative pairs. This approach allows for better utilization of the world knowledge embedded in the pre-trained LVLM, while quickly adapting to the CIR task through straightforward alignment training. To enhance the model’s understanding of both fine-grained and coarse-grained visual information, we propose a module named Pyramid Patcher, inspired by the multi-scale technique in visual detection (Singh and Davis 2018). Unlike the multi-scale technique, which randomly introduces images of different scales during training to adapt to varying object sizes during testing, Pyramid Patcher divides an image into multiple tokens with different visual receptive fields. This approach significantly improves the capability of LVLM without introducing excessive computational overhead. After training the Pyramid Matching Model, we can obtain an initial retrieval result for any composed query.
@@
In supervised CIR tasks, the retrieval model is required to be trained on a training set within a specific domain and tested within it to evaluate performance. Some existing multi-stage methods (Liu et al. 2023b, 2025) additionally train a ranking model to reorder the retrieval results for more accurate results. However, training a ranking model requires reconstructing the training data and consumes additional computational resources. This leads us to consider whether it is possible to refine retrieval results without training. Achieving this without training while providing informational gain often implies a stronger generalization ability. This directs our attention to Chain-of-Thought (Wei et al. 2022), which can enhance the capabilities of models without additional training. In zero-shot CIR tasks (Yang et al. 2024), some methods (Sun, Ye, and Gong 2023; Tang et al. 2025b) incorporate the Chain-of-Thought to generate more accurate textual descriptions of reference images and modified instructions. After that, these methods utilized the CLIP (Radford et al. 2021) model for text and image representation extraction to perform retrieval. Based on these considerations, we designed a Training-Free Refinement paradigm, which adjusts the retrieval results of the Pyramid Matching Model, without the need to train an additional ranking model, thus achieving further performance improvements.
@@
However, directly adopting such Reasoning methods presents two problems: 1) In zero-shot CIR tasks, these methods aim to produce better textual descriptions, while in our Training-Free Refinement stage, we aim to obtain a better refinement score. 2) Existing reasoning-based methods rely on explicit textual reasoning paths, which often lead to significant consumption of computing resources. To address these empirical research gaps, we first pair the composed query with the candidate images retrieved by the Pyramid Matching Model, then utilize the pre-trained LVLM to determine whether they form a correct combination to obtain a refinement score. Inspired by representation engineering (Tang et al. 2025a; Zou et al. 2023), which extracted the representations of LLMs with data that reflects specific capabilities, and treats these representations as the fundamental unit of analysis to understand and control high-level capabilities of LLMs. We designed a reasoning-augmented representation (RAug-Rep) extraction and injection method in the Training-Free Refinement stage. On the training set, we extract the ‘RAug-Rep’ using pre-generated reasoning paths. During the inference phase, we inject the ‘RAug-Rep’ into the intermediate layers of the model to participate in the forward propagation of model. Interestingly, this approach effectively improves the accuracy of the refinement scores, as if finding a key to unlock a certain capability of the model. By simply inserting it into the model, we can stimulate a certain capability of the model, thereby achieving a performance improvement without inputting any explicit textual reasoning paths. Our main contributions are summarized as follows:
@@
Composed Image Retrieval (CIR) (Vo et al. 2019a) is a challenging task due to its reliance on multimodal information for target image retrieval. Most works focus on how to better integrate multimodal information, such as early-fusion (Wen et al. 2024; Levy et al. 2024) and late-fusion (Baldrati et al. 2022; Chen, Zhou, and Peng 2024). Some methods (Liu et al. 2023b, 2025) employ a multi-stage paradigm, where a ranking model is trained in the second stage to further enhance retrieval performance. However, training a ranking model requires reconstructing the training data and consumes additional computational resources. With the rapid development of Chain-of-Thought (Wei et al. 2022), many works (Sun, Ye, and Gong 2023; Tang et al. 2025b; Yang et al. 2024) utilize this technique in zero-shot CIR tasks to generate better text descriptions. Nevertheless, these reasoning-augmented texts depend on complex prompt designs and explicit reasoning paths, posing significant challenges for implementation in embedding-based retrieval tasks like supervised CIR. Based on this, we propose the PMTFR, which combines the advantages of multi-stage and reasoning-augment, achieving better performance with higher efficiency.
@@
Large Vision-Language Models (LVLMs) have made significant strides in multimodal learning, such as BLIP (Li et al. 2022), LLaVA (Liu et al. 2023a), and Qwen-VL (Bai et al. 2023). The LVLMs have achieved excellent results in various multimodal tasks, including visual grounding (Deng et al. 2021) and visual question answering (Antol et al. 2015), among others. In recent years, some CIR methods (Sun et al. 2025; Bai et al. 2024; Li et al. 2025) have started leveraging the capabilities of LVLMs. They use LVLMs to enhance model performance by generating better textual descriptions (Tang et al. 2025b), sentence-level prompting (Bai et al. 2024), or soft prompting (Sun et al. 2025). These works mostly focus on complex designs for text, thereby overlooking the exploration of rich visual cues. In contrast, we leverage the LVLM as an encoder, simply inputting the composite query and target image to obtain the corresponding representations. To enhance the understanding of visual information, we propose a straightforward Pyramid Patcher module, which enables the model to handle information at different visual granularities.
@@
Representation Engineering (RepE)  (Zou et al. 2023) extracted the representations of large language models (LLMs) with data that reflects specific capabilities, and treats these representations as the fundamental unit of analysis to understand and control high-level capabilities of LLMs. As a widely applied technology, it is often used in LLMs and applied in many areas, such as hallucination alleviation (Arditi et al. 2024; Li et al. 2023), instruction following (Stolfo et al. 2024), and reasoning (Tang et al. 2025a). In recent years, there have also been works (Tian et al. 2025a; McGrath et al. 2022) beginning to explore its potential applications in LVLMs. In this work, we pioneer the adaptation of RepE to composed image retrieval (CIR), where reasoning-augmented representations are extracted and injected via a training-free paradigm to boost retrieval performance significantly.
@@
Pyramid Patcher. In a standard Vision Transformer (ViT) (Dosovitskiy et al. 2021), an input 2D image $\bm{I}\in\mathbb{R}^{H\times W\times C}$ is divided into patches, resulting a sequence of flattened 2D patches $\bm{I}^{p}\in\mathbb{R}^{L\times\left(P^{2}\cdot C\right)}$, where $\left(H,W\right)$ is the resolution of the original image, $C$ is the number of channels, $P$ is the patch size and $L=HW/P^{2}$. The final token sequence fed into the visual encoder can be formulated as:
@@
where $f_{\zeta}\left(\cdot\right)$ denotes the Pyramid Patcher. For a sample pair $\langle\bm{u}^{q},\bm{u}^{t}\rangle$, we use the InfoNCE (He et al. 2020) loss to minimize the similarity between the positive pairs and maximize between the negative pairs. Specifically, let $N$ denote the size of the mini-batch and denote $\bm{\Psi}$ the set of all target image representations in a mini-batch, then the loss function can be expressed as:
@@
Reasoning-Augmented Representation Construction. Recent works such as DeepSeek-R1 (Guo et al. 2025) have demonstrated the surprising generalization capabilities of Chain-of-Thought (COT) (Wei et al. 2022) across many tasks. However, this ability relies on explicit textual reasoning, which is not naturally compatible with embedding-based retrieval approaches such as CIR. Fortunately, with the help of representation engineering (Zou et al. 2023), we extract an ‘RAug-Rep’ that benefits from the reasoning ability of the model and enables model performance gains. Specifically, for a positive pair $\langle\bm{Q}_{i},\tilde{\bm{I}}_{i}^{t}\rangle$ in the training set, we first use the ’Refinement Prompt’ (see Appendix A) to construct the question $\bm{q_{i}}$. We then employ the ’Reasoning Path Construction Prompt’ (see Appendix A) to generate input for GPT-4o (Hurst et al. 2024), obtaining the reasoning path denoted as $\bm{c}_{i}$. For each of these inputs, we feed them separately into the pre-trained LVLM to extract layer-wise representations. Formally, for the standalone question $\bm{q_{i}}$ and the augmented input $(\bm{q_{i}};\bm{c}_{i})$ (where ; denotes concatenation), the corresponding representations can be formulated as:
@@
where $\bm{h}_{L}^{0}$ denotes the 0th token of the hidden states in layer $L$ and $\alpha$ is the hyperparameter to balance the strength of injection. After that, following (Liu et al. 2024a), we scaled the magnitude of the vectors back to their original scale:
@@
Following SPRC (Bai et al. 2024), we have selected the two most commonly used benchmarks for the supervised CIR task: Fashion-IQ (Wu et al. 2021) and CIRR (Liu et al. 2021). (1) Fashion-IQ comprises 77,684 fashion images, which are carefully organized into 30,134 triplets and categorized into three distinct subsets: Dress, Shirt, and Toptee. Following previous work (Bai et al. 2024; Sun et al. 2025), we trained our model on the training set and reported the Recall@K metrics for three subsets on the validation set. (2) CIRR is a general image dataset that comprises 36,554 triplets. Following previous work (Bai et al. 2024; Liu et al. 2021), we trained our model in the training set and performed hyperparameter selection on the validation set. Finally, we submitted the prediction results of the test set to an online server to obtain the Recall@K metrics on the entire set and subsets.
@@
We used the pre-trained Qwen2-VL-7B (Bai et al. 2023) model to initialize the weights of the Pyramid Matching Model and employed the AdamW optimizer, with a learning rate set to 4e-5 for 2 epochs. The temperature in InfoNCE was set to 0.005, with a batch size of 1024 for the Fashion-IQ dataset and 512 for the CIRR dataset. N is set to 100 in the selection of the top N candidates. The pre-trained Qwen2-VL-7B is used in the Training-Free Refinement All hyperparameters were selected using the validation set and held constant across experiments. All hyperparameter selections were made on the validation set and applied to all experiments. All experiments were conducted on a single NVIDIA A800 machine.
@@
To evaluate the superiority of PMTFR, we compare our work with established baselines such as TIRG (Vo et al. 2019b), ARTEMIS (Delmas et al. 2022), TG-CIR (Wen et al. 2023), SPRC (Bai et al. 2024),bi (Liu et al. 2024b), CASE (Levy et al. 2024), DWC (Huang et al. 2024) and FashionERN (Chen et al. 2024), multi-stage work including Re-ranking (Liu et al. 2023b), as well as with the latest accepted works, including CIR-LVLM (Sun et al. 2025), Encoder (Li et al. 2025), CCIN (Tian et al. 2025b).
@@
As shown in Fig. 5 (a), we visualized the $\bm{R}_{i,L}^{q}$ (q) and $\bm{R}_{i,L}^{c}$ (q + c) mentioned in Eq. 6 using the t-SNE (Maaten and Hinton 2008) and selected the last several layers for display. Meanwhile, we replaced the reasoning path with noisy text (q + n), and illustrated its distribution in Fig. 5 (b). It can be observed that inputs containing richer informational content (reasoning paths) yield representations with greater distributional divergence from the original. This indicates that the information density of inputs manifests in their representation distributions. Based on these, during the reasoning-augmented representation Construction phase, we used the $\bm{R}_{i,L}^{c}-\bm{R}_{i,L}^{q}$ as the final ‘RAug-Rep’. The aim is to utilize this difference as a medium, allowing the model to acquire approximate capabilities without depending on explicit reasoning paths.

### v2

@@ Introduction
In recent years, Composed Image Retrieval (CIR) (vo2019composing; goenka2022fashionvlp; wen2023target; bai2024sentencelevel) has gained increasing attention. Compared to traditional image retrieval tasks (gordo2016deep; liu2016deepfashion; chen2022deep), CIR requires not only a reference image as input but also a textual description of the expected modifications to the reference image. The complexity of understanding multi-modal data and modeling cross-modal relationships presents technical challenges, but also enables novel applications of CIR methods in fields such as e-commerce (shankar2017deep) and Internet search (deldjoo2023review).
@@
Large Vision-Language Models (LVLMs) have demonstrated promising multimodal understanding capabilities, and have been adopted in various multimodal tasks, including visual grounding (deng2021transvg), visual question answering (antol2015vqa), and CIR tasks (bai2024sentencelevel; sun2025leveraging; karthik2024vision; sun2023training). However, some of these works (karthik2024vision; tang2025reason) use LVLMs to extract better textual descriptions, which are then used to retrieve images, while others (sun2025leveraging; bai2024sentencelevel) rely on complex input prompt designs. These cross-modal generation—such as image captioning—inevitably leads to information loss. For that sake, these methods may fall short in CIR tasks which require both multi-granular visual information and cross-modal correlation understanding. For example, given the modified instruction ‘dog sticks out its tongue on the grass’ (in Fig. 1), the CIR models need to consider both macro-level background information ‘on the grass’ and micro-level details of the dog’s tongue.
@@
Based on these considerations, we propose training a Pyramid Matching Model by utilizing a pre-trained LVLM. By removing the large language model (LLM) head and taking the last token as representation, we simply input composed queries and target images, and use InfoNCE (he2020momentum) loss to minimize the distance between positive pairs while maximizing the distance between negative pairs. This approach allows for better utilization of the world knowledge embedded in the pre-trained LVLM, while quickly adapting to the CIR task through straightforward alignment training. To enhance the model’s understanding of both fine-grained and coarse-grained visual information, we propose a module named Pyramid Patcher, inspired by the multi-scale technique in visual detection (singh2018analysis). Unlike the multi-scale technique, which randomly introduces images of different scales during training to adapt to varying object sizes during testing, Pyramid Patcher divides an image into multiple tokens with different visual receptive fields. This approach significantly improves the capability of LVLM without introducing excessive computational overhead. After training the Pyramid Matching Model, we can obtain an initial retrieval result for any composed query.
@@
In supervised CIR tasks, the retrieval model is required to be trained on a training set within a specific domain and tested within it to evaluate performance. Some existing multi-stage methods (liu2023candidate; liu2025lamra) additionally train a ranking model to reorder the retrieval results for more accurate results. However, training a ranking model requires reconstructing the training data and consumes additional computational resources. This leads us to consider whether it is possible to refine retrieval results without training. Achieving this without training while providing informational gain often implies a stronger generalization ability. This directs our attention to Chain-of-Thought (wei2022chain), which can enhance the capabilities of models without additional training. In zero-shot CIR tasks (yang2024ldre), some methods (sun2023training; tang2025reason) incorporate the Chain-of-Thought to generate more accurate textual descriptions of reference images and modified instructions. After that, these methods utilized the CLIP (radford2021learning) model for text and image representation extraction to perform retrieval. Based on these considerations, we designed a Training-Free Refinement paradigm, which adjusts the retrieval results of the Pyramid Matching Model, without the need to train an additional ranking model, thus achieving further performance improvements.
@@
However, directly adopting such Reasoning methods presents two problems: 1) In zero-shot CIR tasks, these methods aim to produce better textual descriptions, while in our Training-Free Refinement stage, we aim to obtain a better refinement score. 2) Existing reasoning-based methods rely on explicit textual reasoning paths, which often lead to significant consumption of computing resources. To address these empirical research gaps, we first pair the composed query with the candidate images retrieved by the Pyramid Matching Model, then utilize the pre-trained LVLM to determine whether they form a correct combination to obtain a refinement score. Inspired by representation engineering (tang2025unlocking; zou2023representation), which extracted the representations of LLMs with data that reflects specific capabilities, and treats these representations as the fundamental unit of analysis to understand and control high-level capabilities of LLMs. We designed a reasoning-augmented representation (RAug-Rep) extraction and injection method in the Training-Free Refinement stage. On the training set, we extract the ‘RAug-Rep’ using pre-generated reasoning paths. During the inference phase, we inject the ‘RAug-Rep’ into the intermediate layers of the model to participate in the forward propagation of model. Interestingly, this approach effectively improves the accuracy of the refinement scores, as if finding a key to unlock a certain capability of the model. By simply inserting it into the model, we can stimulate a certain capability of the model, thereby achieving a performance improvement without inputting any explicit textual reasoning paths. Our main contributions are summarized as follows:
@@
Composed Image Retrieval (CIR) (vo2019composing) is a challenging task due to its reliance on multimodal information for target image retrieval. Most works focus on how to better integrate multimodal information, such as early-fusion (wen2024simple; levy2024data) and late-fusion (baldrati2022conditioned; chen2024spirit). Some methods (liu2023candidate; liu2025lamra) employ a multi-stage paradigm, where a ranking model is trained in the second stage to further enhance retrieval performance. However, training a ranking model requires reconstructing the training data and consumes additional computational resources. With the rapid development of Chain-of-Thought (wei2022chain), many works (sun2023training; tang2025reason; yang2024ldre) utilize this technique in zero-shot CIR tasks to generate better text descriptions. Nevertheless, these reasoning-augmented texts depend on complex prompt designs and explicit reasoning paths, posing significant challenges for implementation in embedding-based retrieval tasks like supervised CIR. Based on this, we propose the PMTFR, which combines the advantages of multi-stage and reasoning-augment, achieving better performance with higher efficiency.
@@
Large Vision-Language Models (LVLMs) have made significant strides in multimodal learning, such as BLIP (li2022blip), LLaVA (liu2023visual), and Qwen-VL (bai2023qwen). The LVLMs have achieved excellent results in various multimodal tasks, including visual grounding (deng2021transvg) and visual question answering (antol2015vqa), among others. In recent years, some CIR methods (sun2025leveraging; bai2024sentencelevel; li2025encoder) have started leveraging the capabilities of LVLMs. They use LVLMs to enhance model performance by generating better textual descriptions (tang2025reason), sentence-level prompting (bai2024sentencelevel), or soft prompting (sun2025leveraging). These works mostly focus on complex designs for text, thereby overlooking the exploration of rich visual cues. In contrast, we leverage the LVLM as an encoder, simply inputting the composite query and target image to obtain the corresponding representations. To enhance the understanding of visual information, we propose a straightforward Pyramid Patcher module, which enables the model to handle information at different visual granularities.
@@
Representation Engineering (RepE)  (zou2023representation) extracted the representations of large language models (LLMs) with data that reflects specific capabilities, and treats these representations as the fundamental unit of analysis to understand and control high-level capabilities of LLMs. As a widely applied technology, it is often used in LLMs and applied in many areas, such as hallucination alleviation (arditi2024refusal; li2023inference), instruction following (stolfoimproving), and reasoning (tang2025unlocking). In recent years, there have also been works (tian2025representation; mcgrath2022acquisition) beginning to explore its potential applications in LVLMs. In this work, we pioneer the adaptation of RepE to composed image retrieval (CIR), where reasoning-augmented representations are extracted and injected via a training-free paradigm to boost retrieval performance significantly.
@@
Pyramid Patcher. In a standard Vision Transformer (ViT) (dosovitskiyimage), an input 2D image $\bm{I}\in\mathbb{R}^{H\times W\times C}$ is divided into patches, resulting a sequence of flattened 2D patches $\bm{I}^{p}\in\mathbb{R}^{L\times\left(P^{2}\cdot C\right)}$, where $\left(H,W\right)$ is the resolution of the original image, $C$ is the number of channels, $P$ is the patch size and $L=HW/P^{2}$. The final token sequence fed into the visual encoder can be formulated as:
@@
where $f_{\zeta}\left(\cdot\right)$ denotes the Pyramid Patcher. For a sample pair $\langle\bm{u}^{q},\bm{u}^{t}\rangle$, we use the InfoNCE (he2020momentum) loss to minimize the similarity between the positive pairs and maximize between the negative pairs. Specifically, let $N$ denote the size of the mini-batch and denote $\bm{\Psi}$ the set of all target image representations in a mini-batch, then the loss function can be expressed as:
@@
Reasoning-Augmented Representation Construction. Recent works such as DeepSeek-R1 (guo2025deepseek) have demonstrated the surprising generalization capabilities of Chain-of-Thought (COT) (wei2022chain) across many tasks. However, this ability relies on explicit textual reasoning, which is not naturally compatible with embedding-based retrieval approaches such as CIR. Fortunately, with the help of representation engineering (zou2023representation), we extract an ‘RAug-Rep’ that benefits from the reasoning ability of the model and enables model performance gains. Specifically, for a positive pair $\langle\bm{Q}_{i},\tilde{\bm{I}}_{i}^{t}\rangle$ in the training set, we first use the ’Refinement Prompt’ (see Appendix A) to construct the question $\bm{q_{i}}$. We then employ the ’Reasoning Path Construction Prompt’ (see Appendix A) to generate input for GPT-4o (hurst2024gpt), obtaining the reasoning path denoted as $\bm{c}_{i}$. For each of these inputs, we feed them separately into the pre-trained LVLM to extract layer-wise representations. Formally, for the standalone question $\bm{q_{i}}$ and the augmented input $(\bm{q_{i}};\bm{c}_{i})$ (where ; denotes concatenation), the corresponding representations can be formulated as:
@@
where $\bm{h}_{L}^{0}$ denotes the 0th token of the hidden states in layer $L$ and $\alpha$ is the hyperparameter to balance the strength of injection. After that, following (liu2024context), we scaled the magnitude of the vectors back to their original scale:
@@
Following SPRC (bai2024sentencelevel), we have selected the two most commonly used benchmarks for the supervised CIR task: Fashion-IQ (Wu_2021_fashioniq) and CIRR (liu2021CIRR). (1) Fashion-IQ comprises 77,684 fashion images, which are carefully organized into 30,134 triplets and categorized into three distinct subsets: Dress, Shirt, and Toptee. Following previous work (bai2024sentencelevel; sun2025leveraging), we trained our model on the training set and reported the Recall@K metrics for three subsets on the validation set. (2) CIRR is a general image dataset that comprises 36,554 triplets. Following previous work (bai2024sentencelevel; liu2021CIRR), we trained our model in the training set and performed hyperparameter selection on the validation set. Finally, we submitted the prediction results of the test set to an online server to obtain the Recall@K metrics on the entire set and subsets.
@@
We used the pre-trained Qwen2-VL-7B (bai2023qwen) model to initialize the weights of the Pyramid Matching Model and employed the AdamW optimizer, with a learning rate set to 4e-5 for 2 epochs. The temperature in InfoNCE was set to 0.005, with a batch size of 1024 for the Fashion-IQ dataset and 512 for the CIRR dataset. N is set to 100 in the selection of the top N candidates. The pre-trained Qwen2-VL-7B is used in the Training-Free Refinement All hyperparameters were selected using the validation set and held constant across experiments. All hyperparameter selections were made on the validation set and applied to all experiments. All experiments were conducted on a single NVIDIA A800 machine.
@@
To evaluate the superiority of PMTFR, we compare our work with established baselines such as TIRG (vo2019tirg), ARTEMIS (delmasartemis), TG-CIR (wen2023target), SPRC (bai2024sentencelevel),bi (liu2024bi), CASE (levy2024data), DWC (huang2024dynamic) and FashionERN (chen2024fashionern), multi-stage work including Re-ranking (liu2023candidate), as well as with the latest accepted works, including CIR-LVLM (sun2025leveraging), Encoder (li2025encoder), CCIN (tian2025ccin).
@@
As shown in Fig. 5 (a), we visualized the $\bm{R}_{i,L}^{q}$ (q) and $\bm{R}_{i,L}^{c}$ (q + c) mentioned in Eq. 6 using the t-SNE (maaten2008visualizing) and selected the last several layers for display. Meanwhile, we replaced the reasoning path with noisy text (q + n), and illustrated its distribution in Fig. 5 (b). It can be observed that inputs containing richer informational content (reasoning paths) yield representations with greater distributional divergence from the original. This indicates that the information density of inputs manifests in their representation distributions. Based on these, during the reasoning-augmented representation Construction phase, we used the $\bm{R}_{i,L}^{c}-\bm{R}_{i,L}^{q}$ as the final ‘RAug-Rep’. The aim is to utilize this difference as a medium, allowing the model to acquire approximate capabilities without depending on explicit reasoning paths.